OpenStudy (anonymous):

which ordered pair is. a solution y= x-3 A.(-2,5) B.(-5,2) C.(2,5) D.(5,2)

OpenStudy (cggurumanjunath):

substitute each value of a,b,c,d in y=x-3 if lhs=rhs then it is a solution.

OpenStudy (cggurumanjunath):

lhs==left hand side rhs==right hand side say in a x=-2 y=5 y=x-3 lhs=y=5 rhs=x-3=-2-3=-5 here lhs is not equal to rhs similarly try b,c and d.

OpenStudy (anonymous):

listen to him

OpenStudy (anonymous):

What about for y=4x A.(1,3) B.(-1,-4) C.(-4,-1) D.(1,4)

OpenStudy (anonymous):

it is in slope intercept form right now. y=mx+b and b is were it crosses y

OpenStudy (cggurumanjunath):

u try i will help.

OpenStudy (anonymous):

I don't get how to do it I still didn't get the other awnser

OpenStudy (anonymous):

just start from 0 and make a line going up 4 over 1 and same going down